Solve the following question through simplex method
Max Z= 4X1 + 3X2
S+C
2X1 +X2 < 1000
X1 + X2 < 800
X1< 400
X2< 700
X1, X2 >0
1
Expert's answer
2021-09-29T17:13:38-0400

The problem is converted to canonical form by adding slack, surplus and artificial variables as appropriate


1. As the constraint-1 is of type '≤' we should add slack variable "S_1"


2. As the constraint-2 is of type '≤' we should add slack variable "S_2"


3. As the constraint-3 is of type '≤' we should add slack variable "S_3"


4. As the constraint-4 is of type '≤' we should add slack variable "S_4"


After introducing slack variables

"\text{Max } z=4x_1+3x_2+0S_1+0S_2+0S_3+0S_4\\
\text{Subject to }\\
~~2x_1+x_2+S_1~~~~~~~~~~~~~~~~~~~~~~~~=1000\\
~~~~x_1+x_2~~~~~~~~+S_2~~~~~~~~~~~~~~~~=800\\
~~~~x_1~~~~~~~~~~~~~~~~~~~~~~~~~+S_3~~~~~~~~=400\\
~~~~~~~~~~~~~x_2~~~~~~~~~~~~~~~~~~~~~~~~+S_4=700\\
\text{All variables nonnegative}"



Negative minimum "z_j-c_j"  is -4 and its column index is 1. So, the entering variable is "x_1" .


Minimum ratio is 400 and its row index is 3. So, the leaving basis variable is "S_3" .


∴ The pivot element is 1.

"R_3(new)=R_3(old), R_1(new)=R_1(old)-2R_3(new)\\
R_2(new)=R_2(old)-R_3(new), R_4(new)=R_4(old)"


Negative minimum "z_j-c_j"  is -3 and its column index is 2. So, the entering variable is "x_2" .


Minimum ratio is 200 and its row index is 1. So, the leaving basis variable is "S_1" .


∴ The pivot element is 1.

"R_1(new)=R_1(old), R_2(new)=R_2(old)-R_1(new)\\
R_3(new)=R_3(old), R_4(new)=R_4(old)-R_1(new)"



Negative minimum "z_j-c_j"  is -2 and its column index is 5. So, the entering variable is "S_3" .


Minimum ratio is 200 and its row index is 2. So, the leaving basis variable is "S_2" .


∴ The pivot element is 1.

"R_2(new)=R_2(old), R_1(new)=R_1(old)+2R_2(new)\\
R_3(new)=R_3(old)-R_2(new), R_4(new)=R_4(old)-2R_2(new)"


Since all "z_j-c_j\geq0"

Hence, optimal solution is arrived with value of variables as :

"x_1=200,x_2=600"


"Max~z=2600"